### v2

# Entropy-Gated Branching for Efficient Test-Time Reasoning

Abstract
Test-time compute methods like beam search can significantly improve the reasoning capabilities and problem-solving accuracy of large language models. However, these approaches require substantially increased computational resources, with most computation wasted on exploring low-diversity branches where the model already exhibits high confidence. We observe that a small subset of uncertain reasoning steps has a disproportionately large impact on final prediction accuracy, and branching at these points tends to yield higher-quality and more diverse candidate reasoning steps. Therefore, we introduce Entropy‑Gated Branching: a novel inference technique that dynamically allocates computational resources by selectively expanding prediction sequences only at points of high uncertainty. Our method leverages entropy as a gating mechanism to identify when branching is most beneficial, coupled with an external feedback model to rank and prune candidate branches. Empirical results on mathematical and financial reasoning benchmarks show that this strategy improves accuracy by 22.6% over standard inference while operating 37% faster than conventional beam search with similar or higher performance. Our results show that dynamic resource allocation during inference can substantially improve both efficiency and effectiveness, offering a more scalable pathway to enhanced LLM reasoning capabilities.
1 Introduction
Figure 1: Entropy (blue) and Varentropy (red) distribution of Llama-3.2-1B-instruct solving a Chartered Financial Analyst (CFA) problem, demonstrating uncertainty spikes aligning with the model’s mistakes.
Language models have demonstrated remarkable capabilities across a diverse range of tasks (Bang et al. 2023; Li et al. 2023; Mahfouz et al. 2024), yet persist in making fundamental errors due to hallucinations or flawed reasoning (Huang et al. 2024; Tong et al. 2024; Farquhar et al. 2024; Zhang et al. 2025; Galitsky 2025; Penny-Dimri et al. 2025). These mistakes are especially detrimental in high-stakes domains such as finance and healthcare. To address this issue, it is crucial to develop models that ensure outputs are both verifiable and interpretable, reducing the risk of critical mistakes.
Consequently, understanding the underlying causes of these errors and devising practical mitigation strategies is vital for building more reliable language models. Many studies (Xie et al. 2025; Yu et al. 2025; Wei et al. 2025) have demonstrated that fine-tuning language models using rewards or high-quality data can significantly enhance their performance. However, these approaches require substantial time and computational resources to retrain the model. Recently, the research community has shifted focus to test-time compute (TTC) as a complementary approach to improving model performance at low cost (Bi et al. 2024; Liang et al. 2024b; Beeching, Tunstall, and Rush 2024; Snell et al. 2025). TTC methods offer the compelling advantage of enhancing model capabilities without modifying trained weights, enabling performance scaling at inference time. Most TTC methods operationalize extra computation through extensive search algorithms, expanding multiple reasoning branches at every step and selecting the most promising continuation.
However, indiscriminately exploring all candidate branches at every step is very time-consuming. Test-time search algorithms could take minutes to explore all potential intermediate reasoning chains before generating the final answer. This is not user-friendly and poses a significant barrier to real-world deployment. To understand when that extra test-time computation is worthwhile, we analyzed the entropy of the next-token logit distributions on mathematical reasoning benchmarks (Wang et al. 2025). As shown in Figure 1, models are considerably more error-prone during entropy spikes. Since a spike implies a broad posterior over tokens, allocating extra expansion precisely at these high-entropy steps yields maximal diversity and a better chance to rescue an otherwise faulty trajectory; conversely, when entropy is low, the model exhibits high confidence, making further expansion less fruitful since it would likely produce similar continuations. Empirically, we found that 39 out of 50 occurrences that coincide with a high-entropy spike lead to a flawed step of reasoning. These observations suggest that high entropy moments can indicate effective points to branch the model’s reasoning, rather than wasting computation by branching at every step. This raises the question: Is it possible to improve TTC efficiency by selectively expanding beams while still maintaining the performance benefits of beam search?
We propose Entropy-Gated Branching (EGB), a novel TTC inference method that uses the entropy of the model’s predicted logit distribution to identify optimal branching points, instead of expanding at every step. We adopt entropy as a measure of uncertainty to concentrate search budget on high-impact uncertainty moments, while confident beams continue with standard sampling, avoiding unnecessary computational overhead. EGB demonstrates the ability to match or outperform test-time beam search, while achieving substantial reductions in inference time and computational costs.
2 Related Work
Uncertainty Quantification
Handling uncertainty in language models is crucial for ensuring reliable decision-making and response generation. Prior work (Manakul, Liusie, and Gales 2023; Abbasi-Yadkori et al. 2024; Wagner et al. 2024) has explored sampling multiple responses to detect inconsistencies as a measure of uncertainty. However, despite their effectiveness, those methods cannot assess the reliability of a single response, require multiple iterations instead, making them inefficient for real-world applications. An alternative approach leverages token-probability methods (Fadeeva et al. 2024; Duan et al. 2024), which assess uncertainty at the token level without requiring multiple resampling steps.
Test-Time Computation
Recent research (Snell et al. 2025; Liu et al. 2025; Ji et al. 2025) has demonstrated that instead of allocating substantial GPU resources for training new models, enhancing computational resources during inference can yield significant performance gains. Currently, most test-time computing methods rely on generating multiple solutions through repeated sampling and then selecting the best answer. Many search algorithms like Monte-Carlo Tree Search (Zhou et al. 2023; Zhang et al. 2023; Liu et al. 2024a; Koh et al. 2024), guided beam search (Xie et al. 2024) or hybrid search (Snell et al. 2025) show that using search at inference time can lead to performance gains in various tasks. Another line of work (Muennighoff et al. 2025) appends a “wait” token as a method to control the budget and enhance a model’s reasoning process during inference.
LLM Feedback
In language models, feedback is often introduced to improve performance (Pan et al. 2024; Li 2025; Estévez-Ayres et al. 2025). Recent studies have explored various methods to align language models with human preferences (Liang et al. 2024a; Scheurer et al. 2023), modifying prompts (Zhou et al. 2022; Liu et al. 2024b; Ma et al. 2024) or refining errors (Jimichi, Funakoshi, and Okumura 2023; Kirstein, Ruas, and Gipp 2024). Integrating more sophisticated feedback loops will be essential for developing language models that are more reliable to diverse applications.
Figure 2: Illustration of the Entropy-Gated Branching: Left: The normal decoding where the model flows naturally. Middle: Traditional beam search samples KWKW candidates and depend on PRM score keep the top $K$ beams. Right: Entropy-gated branching expands uncertain beams at high entropy moments and generates confident beams normally.
3 Methodology
[…]
The principal limitation of this approach is its rigidly uniform search breadth. Exploring a fixed number of branches at every step often wastes significant compute when the model is confident and unable to produce diverse, high-quality alternate candidate sequences.
3.1 Entropy-Gated Branching
To improve computational efficiency, we propose Entropy-Gated Branching (EGB), a selective expansion strategy that allocates resources only where uncertainty warrants further exploration. Rather than expanding every beam uniformly, EGB uses the entropy of the model’s logit distribution as a dynamic gating mechanism. The entropy of a model’s logit distribution at timestep $t$ is defined as:
$H_{t}=-\sum_{i=1}^{V}p_{i,t}\log_{2}p_{i,t},$ (1)
where $p_{i,t}$ denotes the predicted probability of token $i$ at step $t$, and $V$ is the vocabulary size. Higher entropy corresponds to greater uncertainty and indicates a more uniform probability distribution over tokens.
We also calculate varentropy as shown in Figure 1, measuring the variance in entropy values over multiple potential outputs, defined as:
$V_{t}=\sum_{i=1}^{V}p_{i,t}\left(\log_{2}p_{i,t}+H_{t}\right)^{2}$ (2)
By capturing fluctuations in uncertainty, varentropy highlights situations where the model exhibits inconsistent confidence levels across tokens and helps us get a better understanding of the high-entropy moments.
This leads to two distinct cases for handling beams at each step, based on a pre-defined entropy threshold, $\tau$.
Certain Beams $\mathcal{C}_{t}=\{b\mid H_{t}^{(b)}\leq\tau\}$
When the entropy for a given beam $b$ is below the threshold $\tau$, it signals that the model is confident in its next-token prediction. In this case, EGB avoids wasteful exploration by not creating unnecessary branches for high-confidence parts of the solution. We use the default sampling configuration to generate the most likely next reasoning chunk.
Uncertain Beams $\mathcal{U}_{t}=\{b\mid H_{t}^{(b)}>\tau\}$
When entropy exceeds the threshold, it indicates a critical decision point where the model is uncertain and multiple reasoning paths may be viable. These are the moments where exploration is most valuable. The uncertain beam is expanded into multiple diverse candidate continuations. For a beam $i$, this creates a set of $W$ candidate branches. This concentrates search budget precisely where it is needed most, increasing the likelihood of discovering a correct reasoning path that a normal decoding approach may miss.
3.2 Formalizing Selective Expansion
This process yields a combined pool of candidates, $\mathcal{P}_{t}$, from all active beams. The composition of this pool is determined by whether a beam belongs to the uncertain set $\mathcal{U}_{t}$ or the certain set $\mathcal{C}_{t}$:
$\mathcal{P}_{t}=\bigcup_{i=1}^{K}\begin{cases}\{c_{t,1}^{(i)},\dots,c_{t,W}^{(i)}\},&\text{if }i\in\mathcal{U}_{t}\\ \{c_{t,1}^{(i)}\},&\text{if }i\in\mathcal{C}_{t}\end{cases}$ (3)
For uncertain beams requiring expansion, we employ multinomial sampling with temperature scaling to ensure diversity among the $W$ generated candidates. Specifically, for each uncertain beam $i\in\mathcal{U}t$, we sample from the probability distribution:
$p^{\prime}{j,t}=\frac{\exp(l_{j,t}/\tau_{temp})}{\sum_{k=1}^{V}\exp(l_{k,t}/\tau_{temp})}$ (4)
where $l_{j,t}$ represents the raw logit for token $j$ at timestep $t$, and $\tau_{temp}$ is a temperature parameter that controls the diversity of sampling. A lower temperature encourages more focused sampling around high-probability tokens, while higher temperatures promote greater exploration of the token space.
The total number of candidates generated at step $t$ is therefore at most $|\mathcal{P}_{t}|=K+(W-1)|\mathcal{U}_{t}|$. Since $|\mathcal{U}_{t}|\leq K$, this is always less than or equal to the $KW$ candidates produced by standard beam search. The computational complexity of selective expansion scales as $O(|\mathcal{U}_{t}|\cdot W\cdot V)$ for candidate generation, compared to $O(K\cdot W\cdot V)$ for uniform beam search. In practice, since $|\mathcal{U}_{t}|$ is typically much smaller than $K$ due to the selective nature of branching, EGB achieves substantial computational savings while maintaining search effectiveness.
3.3 Beam Evaluation
Before proceeding to quality assessment, EGB implements a duplicate detection mechanism to avoid wasting computational resources on scoring identical continuations. Since both uncertain beams and certain beams contribute to the candidate pool $\mathcal{P}_{t}$, identical or near-identical sequences can emerge from different sources, particularly when the model exhibits strong preferences for specific token sequences. Therefore, we employ string matching on the token-level and filter out the same branches first and gather the rest diverse branches for further evaluation. After collecting candidate continuations, assessing their quality is critical, as it directly influences the overall performance of the base model. A strong feedback model ensures that the generation follows a coherent and logical trajectory while minimizing errors. We employ a Process Reward Model that evaluates the quality of reasoning steps rather than just final outcomes. The PRM assigns scores to partial solutions based on their logical consistency, mathematical accuracy, and adherence to sound reasoning principles. For a candidate continuation $c_{t,j}^{(i)}$, the PRM score is computed as:
$s_{PRM}(c_{t,j}^{(i)})=f_{PRM}(\text{context},c_{t,j}^{(i)})$ (5)
where $f_{PRM}$ represents the trained process reward model that takes the problem context and the candidate continuation as input. Finally, we sort the scored candidates in descending order and select the top-$K$ highest-scoring beams as the active beams for continued branching in the next step.
4 Experiments
Datasets
CFA exams are renowned for their rigorous assessment of financial, quantitative, and analytical reasoning, making them an excellent proxy for evaluating model performance in high-stakes, real-world financial decision-making scenarios. Mistakes in the financial questions can lead to significant losses, reinforcing the need for precise and reliable reasoning. As official CFA exam questions are not publicly accessible, this study utilizes CFA mock exams purchased from AnalystPrep (AnalystPrep 2024), covering two levels of the CFA program. The dataset comprises multiple-choice and essay questions, each supplemented with corresponding answers, explanations, grading criteria, and metadata indicating the CFA topic associated with each question.
In addition to CFA data, we evaluate our methods on competition and general mathematical reasoning benchmarks of varying difficulty. First, we include an AIME evaluation set drawn from the AI-MO dataset (Art of Problem Solving Community 2025). This corpus contains all 90 problems from the 2022–2024 American Invitational Mathematics Examination (AIME I & II each year) (Art of Problem Solving Community 2025). Each AIME test is widely viewed as hard in difficulty between grade-school word problems and full Olympiad-style proofs (Mathematical Association of America 2025; Wikipedia contributors 2025).
We also assess performance on the MATH benchmark (Hendrycks et al. 2021), which consists of competition-level problems (AMC, AIME, and other sources) spanning a wide range of high-school mathematics topics and difficulty levels; for all experiments, we use 500 held-out test questions. Finally, we report results on GSM8K (Cobbe et al. 2021), a widely used grade-school math word-problem benchmark designed to probe multi-step arithmetic reasoning in language models, utilizing 1,319 test questions in our experiments.
Language Models
We implement our method on two families of models, Qwen3 (Yang et al. 2025) and Llama3 (Touvron et al. 2023), at 3 sizes each. For Qwen3 we use Qwen3-1.7B, Qwen3-4B, and Qwen3-8B, all in ”non-thinking” mode (Yang et al. 2025). For Llama3 we use Llama-3.2-1B-instruct, Llama-3.2-3B-instruct, and Llama-3.1-8B-instruct. We adopt the recommended sampling configuration for all experiments, with a temperature of 0.7, top-p of 0.8, top-k of 20, and min-p of 0 for Qwen3 with non-thinking mode and a temperature of 0.6, top-p of 0.9 for Llama models
Feedback Models
To provide reliable feedback, we use Qwen2.5-Math-PRM-7B, which is the most popular PRM model on HuggingFace up to date. This model is specifically trained to assess reasoning quality by assigning a numerical score that reflects the correctness of the inference. Unlike the instruct model, which relies on a qualitative analysis to provide a branch selection, this model directly provides a probabilistic evaluation, where a designated probability serves as a quantitative score for ranking the branches. Higher scores indicate stronger logical consistency and correctness.
Branching Settings
To balance performance with computational overhead, we set the entropy thresholds $\tau=2$ for AIME, CFA Level I and Level II math datasets, and $\tau=1.5$ for the GSM8K and MATH500 datasets since models exhibit less uncertainty on general math questions. For results reported in Table 1, we use a beam size $K$ of 3, a beam width $W$ of 5 for all datasets. Detailed analyses of the threshold values and branch numbers can be found in Section 5.3.
Infrastructure
All experiments were performed on A100 or H100 GPUs with 80GB VRAM, on Ubuntu 22, CUDA 12.4, Python 3.12. Each experiment was run on one GPU, utilizing 20-70 GB of GPU memory, depending on the model size and sequence length.
Model Method CFA AIME MATH-500 GSM8K
Level I Level II
Qwen3-1.7B Standard 48.78% 39.77% 6.67% 70.40% 76.80%
Beam Search 54.83% 44.32% 11.11% 75.20% 84.00%
EGB 55.78% 43.18% 11.11% 75.00% 86.35%
Qwen3-4B Standard 67.28% 54.55% 17.78% 82.80% 90.90%
Beam Search 69.11% 56.57% 21.11% 84.00% 89.16%
EGB 70.89% 55.81% 18.60% 84.40% 90.67%
Qwen3-8B Standard 73.50% 51.70% 18.89% 82.40% 92.72%
Beam Search 76.12% 58.86% 18.89% 84.60% 88.25%
EGB 77.22% 58.52% 22.22% 82.20% 92.19%
Llama-3.2-1B-instruct Standard 30.44% 25.57% 2.22% 21.80% 36.39%
Beam Search 39.61% 42.58% 2.22% 39.40% 61.26%
EGB 34.09% 43.91% 2.22% 36.40% 54.89%
Llama-3.2-3B-instruct Standard 43.17% 36.36% 5.56% 39.00% 72.33%
Beam Search 49.06% 44.66% 6.67% 48.20% 82.18%
EGB 50.67% 45.45% 8.89% 51.40% 83.09%
Llama-3.1-8B-instruct Standard 49.17% 38.07% 4.44% 41.80% 83.09%
Beam Search 54.06% 47.73% 6.67% 47.60% 88.74%
EGB 57.50% 49.43% 8.89% 52.40% 88.10%
Table 1: Performance of EGB compared with beam search and standard decoding on finance and math benchmarks.
LLM-based Evaluation
Final answer correctness was determined using a powerful external language model, Gemini 2.5 Flash, as an impartial judge (Gemini Team et al. 2024). This approach circumvents the limitations of rigid evaluation methods like exact string matching or regex-based parsers, which are often brittle and can penalize semantically correct answers due to minor stylistic or formatting variations. For each problem, the LLM judge was provided with the original question, the model’s complete generated response, and the ground-truth solution, and was then prompted to assess whether the final answer was correct. This method ensures a robust and fair assessment by focusing on the semantic correctness of the reasoning outcome rather than superficial syntactic differences, providing a more reliable measure of the model’s true problem-solving capabilities.
5 Results
Our main results are presented in Table 1, outlining the performance of each model across the five benchmark datasets. EGB demonstrates substantial improvements over standard base inference, achieving an average gain of 22.6% across all models and benchmarks with absolute accuracy gains of 4-10 percentage points (pp) in most settings. The method consistently delivers competitive or superior performance compared to beam search while achieving these results with significantly reduced computational cost and time, as detailed in Section 5.2. Our analysis reveals interesting patterns across different model families, scales, and reasoning domains.
5.1 Main results
General Math vs. Financial Reasoning
EGB consistently outperforms standard decoding on all domains, but there are nuanced patterns between models and domains. For smaller models (1-2B parameters), EGB yields comparable improvements across both domains, suggesting that when base model capabilities are limited, both financial and mathematical reasoning benefit equally from increased exploration. However, as model scale increases, a clear trend emerges: financial reasoning tasks maintain robust gains from EGB while general mathematical problems show diminishing returns. At the largest scales tested (8B parameters), financial tasks benefit at approximately twice the rate of general mathematical problems.
This scale-dependent pattern likely reflects fundamental differences in how uncertainty manifests across these reasoning domains. Financial reasoning problems typically require integrative processing that combines numerical computation, contextual interpretation, and domain-specific procedural knowledge. This multi-faceted nature creates what we term a distributed uncertainty landscape, where multiple decision points throughout the reasoning chain contribute meaningfully to solution quality. In such scenarios, EGB’s exploration strategy can effectively navigate the complex interaction between different reasoning components.
Model Families
The Qwen models establish strong baseline performance, with even the smallest Qwen3-1.7B model substantially outperforming comparable Llama models across all benchmarks. For instance, Qwen3-1.7B achieves 48.78% on CFA Level I compared to Llama-3.2-1B’s 30.44%, and 70.40% on MATH-500 versus 21.80% for Llama. This performance advantage persists consistently across all model scales. However, the Llama family demonstrates exceptional responsiveness to EGB, achieving an average improvement of 35.3% over standard decoding compared to Qwen’s more modest 9.9% gains. The Llama models show particularly dramatic improvements on certain tasks. For example, Llama-3.2-1B-instruct improves by +18.34pp on CFA Level II, while the stronger baseline Qwen3-1.7B only gains +3.41pp on the same task. Remarkably, EGB’s substantial gains enable Llama models to bridge much of the performance gap with their Qwen counterparts, often matching or approaching Qwen’s performance despite starting from significantly weaker foundations.
Model Size Trends
EGB’s advantages over beam search increase monotonically with model size. Small models (1-1.7B) show substantial relative gains from EGB (+31.0% over baseline) but remain slightly behind beam search (-1.2pp average). Medium models (3-4B) achieve balanced performance with EGB winning 60% of comparisons against beam search (+2.9% relative advantage). Large models (8B) demonstrate EGB’s strongest performance profile, with a 70% win rate and substantial improvements over beam search (+7.3% relative advantage). This scaling pattern suggests that larger models don’t reduce exploration headroom but rather enhance the quality and effectiveness of exploration itself. The superior performance of larger models with EGB indicates that increased reasoning capability amplifies the benefits of strategic exploration, as these models can both generate more sophisticated alternative approaches and more accurately evaluate their relative merits. While small models benefit from any additional reasoning paths, larger models can more effectively generate diverse, high-quality alternatives and better distinguish between promising and unpromising directions.
Trends in Benchmark Difficulty
EGB shows the greatest impact on tasks where models already have the necessary knowledge but struggle with consistent execution. On benchmarks like GSM8K and the CFA exams, where models show solid baseline competence, EGB provides substantial gains by helping correct reasoning errors and select among viable solution approaches. Even on highly challenging benchmarks like AIME, where baseline scores rarely exceed 22%, EGB drives considerable improvements, pushing models closer to their performance limits. While exploration is naturally constrained when problems go beyond a model’s knowledge, EGB reliably helps models reach their full potential by testing a wide range of reasoning strategies. This allows them to identify and apply their best possible solutions, even at the edge of their capabilities.
The difficulty-correlated performance pattern illuminates EGB’s primary mechanism as systematic optimization of model capabilities rather than knowledge augmentation. On moderate-difficulty tasks, models typically understand problem requirements and possess appropriate solution strategies but may fail due to computational errors or suboptimal path selection. EGB’s diverse reasoning paths provide multiple opportunities to avoid these execution failures. On the most challenging problems, EGB ensures models leverage every available reasoning strategy, maximizing their performance within existing capability constraints.
Figure 3: Average run-time comparison of various inference methods on five benchmarks
5.2 Computational Efficiency Analysis
Run-Time Efficiency
Beyond accuracy improvements, EGB delivers substantial computational efficiency advantages over beam search, achieving 31-75% speedup across all evaluated benchmarks, as shown in Figure 3, while maintaining superior or competitive performance. This efficiency gain represents a significant practical advantage that enhances EGB’s value proposition for real-world deployment scenarios. Unlike traditional accuracy-efficiency trade-offs, EGB achieves both improved performance and reduced computational cost simultaneously. For financial reasoning tasks, EGB delivers 33-40% speedups alongside 2-4 percentage point accuracy improvements. Even on challenging mathematical benchmarks where accuracy gains are modest, the substantial time savings provide clear value for time-sensitive deployments.
GPU Memory Efficiency
EGB provides significant memory efficiency advantages by avoiding the storage overhead associated with maintaining large numbers of parallel candidate sequences. Since most generation steps involve primarily confident predictions, EGB consistently operates with significantly fewer active candidates than traditional beam search, directly translating to proportional memory savings for intermediate state storage, attention caches, and token embeddings. The memory savings compound over longer sequences, where the cumulative effect of selective branching becomes increasingly pronounced.
5.3 Ablation Studies
Figure 4: Impact of beam sizes $K$ and beam widths $W$ on model’s performance, both numbers are scaled from 2 up to 16
Threshold Sensitivity
We explore how the entropy threshold affects model performance on different scales. By analyzing the effect of altering the entropy threshold $H_{\text{thresh}}$ from 1.0 to 4.0 on a subset of 100 questions from MATH500 dataset, Figure 5 shows that $H_{\text{thresh}}=2$ is, on average, the best setting for the Llama models, where performance remains relatively stable within the range of 1.5-3.5. This optimal range highlights that EGB’s effectiveness depends critically on maintaining sufficient branching opportunities during reasoning while avoiding excessive computational overhead. At very low thresholds ($H_{\text{thresh}}<1.5$), the model branches too frequently, often at decision points where the uncertainty is minimal and exploration provides little benefit. This aggressive branching not only increases computational costs but can also introduce noise by exploring fundamentally similar reasoning paths that differ only in minor details or presentation.
Figure 5: Entropy threshold sensitivity analysis using Llama models on MATH-500 subset (20% questions)
Beam Size $K$
We analyze beam size from KK=2 to K=16, revealing that larger beam sizes consistently improve performance, with the most substantial gains occurring between KK=2 and KK=8 across all model scales. The improvement is particularly pronounced for larger models, with Llama-3.1-8B showing dramatic gains from 36.2% to 48.8% accuracy as beam size increases. This pattern confirms that maintaining multiple active reasoning branches helps preserve solution diversity and prevents premature convergence to suboptimal paths. However, the performance plateaus at KK=16, suggesting diminishing returns. This is likely due to a trade-off in PRM-based ranking: as the PRM may favour clusters of similar partial solutions, additional beams may not contribute new information, limiting the benefits of further expansion.
Beam Width $W$
The beam width parameter, controlling how many new candidates are generated at each branching point, shows more modest but consistent improvements as W increases from 2 to 16. While the effect is less dramatic than beam size, larger models still benefit meaningfully from increased exploration breadth, with performance gains of 1-2 percentage points across the range. The relatively smaller impact of beam width compared to beam size suggests that maintaining diverse active branches is more critical than generating many candidates at each branching point. This indicates that EGB’s effectiveness depends more on preserving multiple reasoning trajectories over time than on exhaustive local exploration at individual decision points.
6 Conclusion
We introduced Entropy-Gated Branching (EGB), a selective expansion strategy that dynamically allocates computational resources based on model uncertainty during generation. We addressed the limitation in current test-time compute methods by recognizing that not all generation steps require equal computational investment. By using entropy as a gating mechanism to trigger branching only at high-uncertainty decision points, EGB achieves a rare combination of improved accuracy and reduced computational cost compared to traditional beam search. Our comprehensive evaluation across mathematical reasoning benchmarks demonstrates that EGB consistently delivers 31-75% speedup while maintaining superior or competitive performance to traditional beam search. The method proves particularly effective on moderately challenging tasks where models possess the requisite knowledge but struggle with execution consistency, achieving substantial accuracy improvements on financial and mathematical reasoning problems. By focusing compute on where it matters the most, EGB offers a pragmatic path to robust, cost-aware test-time reasoning at scale.
7 Disclaimer
This work was supported in part by the J.P. Morgan Chase Faculty Research Award, which was prepared for informational purposes in part by the Artificial Intelligence Research Group of JPMorgan Chase & Co. and its affiliates (“JPMorgan”) and is not a product of the Research Department of JPMorgan. JPMorgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy, or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product, or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.